OpenStudy (anonymous):

Solve by completing the square: m^2-49/2m=1/2

OpenStudy (anonymous):

\[m^{2}-49/2m=1/2\]

OpenStudy (anonymous):

m=-1.48

OpenStudy (saifoo.khan):

Hi there. do you want the answer in radicals?

OpenStudy (anonymous):

Hey saifoo, yes please.

OpenStudy (saifoo.khan):

\[\Large \frac{2}{-49-\sqrt{2409}},\frac{1}{4}(49+\sqrt{2409})\]

OpenStudy (anonymous):

m^2 - ((49/2)/2)^2m = 1/2 so [(49/2)2] ^2= 2401/16 add it to both sides m^2 - 49/2m + 2401/16 = 1/2 + 2401/16 (m - 49/4)^2 = 2409/16 m - 49/4 = sqrt(2409)/4 m = [sqrt(2409)/4] +49/4 ---->answer

OpenStudy (anonymous):

hmm new I was off. I got \[\sqrt{2409}/4+7/4 for the first and tottally lost on the second!\]